OpenStudy (anonymous):

Write an equation of a line in slope intercept for that is PARALLEL to y = 3x+6 and passes through the point (-10,2.5) I finally understand slope intercept form until this threw me off! Please help!

OpenStudy (luigi0210):

It's the almost the same thing, the only difference is they already give you the slope. Parallel meaning the same as that equation.

OpenStudy (anonymous):

Parallel lines have the same slope.. y = 3x + 6 So the slope here is 3.. We can plug the slope and the point into point-slope form: \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope. So we have: \(y - 2.5 = 3(x +10)\)

OpenStudy (anonymous):

Simplify and you'll get your answer in slope-intercept form. Start by distributing 3 into the parenthesis.

OpenStudy (anonymous):

thank you! @iGreen.

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

so would I do the same thing for a question that is perpendicular?

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

@DuckDynastyfan923

OpenStudy (luigi0210):

Perpendicular means it's the negative reciprocal of the original so no.

OpenStudy (anonymous):

^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^6